OpenStudy (anonymous):

Find the mass of the lamina described by the inequalities given that its density is p(x,y)=xy. x>= 0, 3 <= y <= 3+(9-x^2)^square root

OpenStudy (anonymous):

The problem also said that polar coordinates could be used to simplify. I just need someone to check to make sure I'm getting the correct limits, as I'm converting..

OpenStudy (anonymous):

\[x \ge0,3 \le y \le 3+\sqrt{9-x^2}\] ^ this is the original problem

OpenStudy (anonymous):

\[\int\limits_{0}^{x}\int\limits_{3}^{3+\sqrt{9-x^2}}(xy)dydx\]

OpenStudy (anonymous):

I know by converting to polar coords one of my r limits is 6 and the other 3, would theta be 0 and pi/2?

OpenStudy (anonymous):

 this is the graph I got for 3+\[\sqrt{9-x^2}\]